# Tight concentration inequalities for quantum adversarial setups exploiting permutation symmetry

### v2

@@ Abstract
I Introduction and the main results
I.1 General introduction
@@
A concentration inequality is a bound on the probability of obtaining atypical events, i.e., events that are far from the expectation. It is essential in lots of information-theoretic primitives such as cryptography and error correction since it gives a failure probability in anticipating that the observed values are close to the expectation values. The same applies to quantum information theory; we need concentration inequalities to obtain an upper bound on the probability of atypical events in quantum information protocols. Evaluation of the probability of atypical events may be easy for independent and identically distributed (i.i.d.) quantum states. There, one can apply Chernoff-Hoeffding theorem [1] or Sanov’s theorem [2, 3] to obtain an upper bound on the deviation from the expectation value for an additive observable allowing a small failure probability. For certain applications, however, one even needs a concentration inequality that applies to an adversarial setup. This includes quantum key distribution (QKD), blind quantum computation, and the verification of a quantum state prepared by an adversary, where the underlying statistics may be governed by an adversary.
@@
In the field of QKD, for example, proving the security against collective attacks in a finite number of communication rounds is relatively easy, and many QKD protocols have been proved secure at least against collective attacks even in the finite-size regime. The situation gets involved when considering the security against general attacks. There are two mainstreams in the way of proving the security against general attacks in finite-size regimes. One of them is called the phase-error correction method [4, 5, 6, 7], which is the main focus of this paper, and evaluates the number of phase error patterns or a phase error rate while allowing an exponentially small failure probability. Correcting the phase error with a failure probability $\epsilon$ ensures $\sim\epsilon^{1/2}$-secrecy of the protocol against the adversary. In Ref. [8], it was shown that in the phase-error correction method, imposing a permutation invariance to the protocol either actively or passively reduces the problem of general attacks to the collective attacks at the cost of increasing the failure probability $\epsilon$ by a polynomial factor. This implies that the error exponent of the failure probability of the phase error correction is the same against collective attacks and against general attacks. After that, a distinct approach using Azuma’s inequality, which is based on sequentiality of measurements instead of permutation invariance, was proposed [9]. Since Azuma’s inequality is easier to evaluate, it prevails in the QKD community. At the cost of its simplicity, the error exponent of the failure probability when using Azuma’s inequality is always less than that of the collective attack case. Intensive research has been made on tightening Azuma’s inequality with additional information [10, 11, 12] including the recently found one with unconfirmed knowledge [12], but it is still looser than the Sanov-type inequalities [2, 3] that apply to the i.i.d. random variables. Note that in another approach to the security proofs based on the leftover hash lemma against a quantum adversary [13], a reduction of the security against general attacks to that against collective attacks by exploiting the permutation symmetry has also been studied under the name of de Finetti reduction [14, 15, 16, 17, 13, 18, 19]. The problems with these approaches were that the reduction was not very straightforward or that they led to worse performance. Then, the community gradually shifted to an alternative method that uses more recently developed entropy accumulation theorem [20, 21, 22]. Interestingly, both Azuma’s inequality and the entropy accumulation utilize sequential conditionings by previous measurement outcomes, yet they may be looser than the bound that applies to the i.i.d. case.
@@
In this paper, we revisit the idea based on permutation invariance and derive concentration inequalities through the reduction to the i.i.d. scenario. The first case we consider is when a quantum state on a multi-qudit system is invariant under the permutation of the subsystems possibly with additional symmetry on each subsystem. We show that whenever an outcome of the quantum measurement has a small appearance probability for any i.i.d. quantum state, it also has a small appearance probability for these permutation-invariant states up to an explicit factor polynomial in the number of subsystems. This can be proved by observing that any permutation-invariant quantum state is bounded from above by the mixture of i.i.d. quantum states up to a polynomial factor. Compared to the known results [13, 18], we tightly evaluate the polynomial prefactor and improve the bound for a general permutation-invariant state. We also obtain a bound that explicitly depends on the irreducible representation space of the permutation group. In particular, when restricted to the totally symmetric subspace, our bound reproduces the one developed for the post-selection technique [13, 18]. The second case we consider is when we perform independent and identical quantum measurements on an arbitrary quantum state prepared by an adversary. In this case, we have an inequality between the probability mass function for an empirical probability of the measurement outcomes and a mixture of multinomial distribution up to a polynomial factor. We then obtain a Sanov-type concentration inequality for this case. As opposed to the first case, the polynomial factor in this case does not depend on the dimension of the underlying quantum system but depends on the number of measurement outcomes.
@@
I.2 Summary of our main results
@@
$Z_{n}^{d}\coloneqq\biggl{\{}\bm{n}=(n_{1},\ldots,n_{d})\in\{0,\ldots,n\}^{\times d}:\sum_{i=1}^{d}n_{i}=n\biggr{\}},$ (3)
@@
$\mathrm{supp}({\cal S}_{A^{n}}(P_{\bm{n}}\rho_{A^{n}}P_{\bm{n}}))$
$\leq f(n,d;\bm{n})\,{\cal S}_{A^{n}}\!\Bigl{(}\mathbb{E}_{U_{A}\sim{\rm Haar}(d)}\Bigl{[}\Bigl{(}U_{A}\,\mathrm{diag}\Bigl{(}\frac{\bm{n}}{n}\Bigr{)}\,U_{A}^{\dagger}\Bigr{)}^{\otimes n}\Bigr{]}\Bigr{)},$ (17)
@@
where we used the fact that $\tau_{A}^{\otimes n}$ is permutation invariant and thus satisfies ${\cal R}_{A^{n}}\circ{\cal S}_{A^{n}}(\tau_{A}^{\otimes n})=\tau_{A}^{\otimes n}$. We see that Eq. (21) takes a similar form to Eq. (1) but with different coefficients $f_{q}(n,d;\bm{n})$. For comparison, let us first replace $f_{q}(n,d;\bm{n})$ with an $\bm{n}$-independent coefficient $f_{q}(n,d)$ using Lemma 1. When $n\gg d$, which is of our interest, the function $f_{q}(n,d)$ scales as $n^{(d^{2}-1)/2}$ whereas the coefficient in the conventional bound (1) scales as $n^{d^{2}-1}$. Our bound thus achieves the square-root improvement in this case.
@@
$\begin{split}&\leq f_{q}(n,d;\bm{n})f_{q}(m,d^{\prime};\bm{m})\\ &\qquad\times{\cal S}_{A^{n}}\!\biggl{(}\mathbb{E}_{U_{A}\sim{\rm Haar}(d)}\Bigl{[}\Bigl{(}U_{A}\,\mathrm{diag}\Bigl{(}\frac{\bm{n}}{n}\Bigr{)}\,U_{A}^{\dagger}\Bigr{)}^{\otimes n}\Bigr{]}\biggr{)}\\ &\qquad\quad\otimes{\cal S}_{B^{m}}\!\biggl{(}\mathbb{E}_{U_{B}\sim{\rm Haar}(d^{\prime})}\Bigl{[}\Bigl{(}U_{B}\,\mathrm{diag}\Bigl{(}\frac{\bm{m}}{m}\Bigr{)}\,U_{B}^{\dagger}\Bigr{)}^{\otimes m}\Bigr{]}\biggr{)},\end{split}$ (26)
@@
${\cal H}_{\bm{j}_{0}(\bm{m})}^{R}\otimes{\cal H}_{\bm{j}_{0}(\bm{m})}^{D}$ $=\bigotimes_{j=1}^{k}\bigl{(}{\cal H}_{j}^{R}\otimes{\cal H}_{j}^{D}\bigr{)}^{\otimes n_{j}}$ (38)
@@
=${\cal R}^{\prime}=\left\{\bm{p}:\bm{u}\cdot\bm{p}\leq\frac{\sqrt{2}\|\bm{u}-(\bm{v}\cdot\bm{u})\bm{v}\|}{n}\right\},$ (55)
@@
Consider the measurement ${\cal M}^{\otimes n}$ on $n$ systems $A^{n}$, which has the POVM $\{M^{\otimes n}(\bm{i})\}_{\bm{i}\in{\cal X}^{\times n}}$ with $M^{\otimes n}\bigl{(}(i_{1},\ldots,i_{n})\bigr{)}=M(i_{1})\otimes\cdots\otimes M(i_{n})$. Let ${\cal A}\subseteq{\cal P}$ be a closed convex region of the simplex. Then, for any quantum state $\rho\in{\cal D}({\cal H}^{\otimes n})$, the probability mass function ${\cal M}^{\otimes n}(\rho):\bm{i}\mapsto\mathrm{Tr}[M^{\otimes n}(\bm{i})\rho]$ satisfies
$\begin{split}\sum_{\bm{i}:\bm{\chi}(\bm{i})/n\in{\cal A}}{\cal M}^{\otimes n}(\rho)(\bm{i})\leq\max_{\bm{p}\in{\cal A},\bm{q}\in{\cal R}^{\prime}}f_{c}(n,k)\exp[-nD(\bm{p}\|\bm{q})],\end{split}$ (56)
@@
Figure 1: Geometrical relations between ${\cal Q}$, ${\cal R}$, ${\cal R}^{\prime}$, and ${\cal A}$ inside the set ${\cal P}$. The blue line is the hyperplane $R\coloneqq\{\bm{p}\in{\cal P}:\bm{u}\cdot\bm{p}=0\}$, which is a boundary of ${\cal R}$. This boundary is shifted by $\sqrt{2}/n$ to form the hyperplane $R^{\prime}$ (the green line), which is a boundary of ${\cal R}^{\prime}$.
@@
I.3 Organization of the paper
In Sec. II, we prove Theorems 13 stated above, where each subsection is assigned for each theorem. In Sec. III, we apply our results to simple examples and numerically demonstrate the tightness of our bounds by comparison with conventionally used bounds. The first example in Sec. III.1 is an estimation task of measurement outcomes with two non-orthogonal projections, which is not directly related to real applications but is still informative. The second example is an application to the quantum key distribution protocol. Finally, in Sec. IV, we summarize our results and discuss possible applications.
II Proofs of the main theorems
II.1 A permutation-invariant quantum state is bounded by a mixture of i.i.d. quantum states
@@ II Proofs of the main theorems
$\begin{split}{\cal S}_{A^{n}}(\tau_{A^{n}})&=\bigoplus_{\bm{n}\in Y_{n}^{d}}\bigl{(}\mathrm{Id}_{{\cal U}_{\bm{n}}}\otimes\mathrm{Tr}_{{\cal V}_{\bm{n}}}(P_{\bm{n}}\tau_{A^{n}}P_{\bm{n}}|_{{\cal U}_{\bm{n}}\otimes{\cal V}_{\bm{n}}})\bigr{)}\\ &\hskip 113.81102pt\otimes\ket{0}\!\bra{0}_{{\cal V}_{\bm{n}}},\end{split}$ (58)
where |$\ket{0}\!\bra{0}_{{\cal V}_{\bm{n}}}$ is an arbitrary pure state on ${\cal V}_{\bm{n}}$. We also define another CPTP map ${\cal R}_{A^{n}}$ as follows:
$\begin{split}{\cal R}_{A^{n}}(\tau_{A^{n}})&=\bigoplus_{\bm{n}\in Y_{n}^{d}}\bigl{(}\mathrm{Id}_{{\cal U}_{\bm{n}}}\otimes\mathrm{Tr}_{{\cal V}_{\bm{n}}}(P_{\bm{n}}\tau_{A^{n}}P_{\bm{n}}|_{{\cal U}_{\bm{n}}\otimes{\cal V}_{\bm{n}}})\bigr{)}\\ &\hskip 113.81102pt\otimes\frac{I_{{\cal V}_{\bm{n}}}}{\dim{\cal V}_{\bm{n}}}.\end{split}$ (59)
@@
$\begin{split}&\mathrm{supp}({\cal S}_{A^{n}}(P_{\bm{n}}\rho_{A^{n}}P_{\bm{n}}))\\ &=\begin{cases}\mathrm{supp}(\rho_{{\cal U}_{\bm{n}}})\otimes\ket{0}\!\bra{0}_{{\cal V}_{\bm{n}}}&\text{on }{\cal U}_{\bm{n}}\otimes{\cal V}_{\bm{n}}\\ 0&\text{otherwise}\end{cases}.\end{split}$ (67)
@@
$=\frac{1}{\alpha_{\bm{n}}(\bm{t})\dim{\cal V}_{\bm{n}}}{\cal S}_{A^{n}}(P_{\bm{n}}\mathbb{E}_{U_{A}\sim\text{Haar}(d)}[(U_{A}\mathrm{diag}(\bm{t})U_{A}^{\dagger})^{\otimes n}]P_{\bm{n}})$ (69)
$\leq\frac{\dim{\cal U}_{\bm{n}}}{s_{\bm{n}}(\bm{t})\dim{\cal V}_{\bm{n}}}{\cal S}_{A^{n}}(\mathbb{E}_{U_{A}\sim\text{Haar}(d)}[(U_{A}\mathrm{diag}(\bm{t})U_{A}^{\dagger})^{\otimes n}]).$ (70)
@@
$\begin{split}&\mathrm{supp}({\cal S}_{A^{n}}(P_{\bm{n}}\rho_{A^{n}}P_{\bm{n}}))\\ &\leq\min_{\bm{t}=(t_{1},\ldots,t_{d})\geq 0:\sum_{i=1}^{d}t_{i}=1}\frac{\dim{\cal U}_{\bm{n}}}{s_{\bm{n}}(\bm{t})\dim{\cal V}_{\bm{n}}}\\ &\qquad\quad{\cal S}_{A^{n}}\!\Bigl{(}\mathbb{E}_{U_{A}\sim{\rm Haar}(d)}\Bigl{[}\bigl{(}U_{A}\,\mathrm{diag}(\bm{t})\,U_{A}^{\dagger}\bigr{)}^{\otimes n}\Bigr{]}\Bigr{)}\end{split}$ (73)
@@
II.2 Generalization to the case with symmetry restrictions

### v3

@@ Abstract
1 Introduction and the main results
1.1 General introduction
@@
A concentration inequality is a bound on the probability of obtaining atypical events, i.e., events that are far from the expectation. It is essential in lots of information-theoretic primitives such as cryptography and error correction since it gives a failure probability in anticipating that the observed values are close to the expectation values. The same applies to quantum information theory; we need concentration inequalities to obtain an upper bound on the probability of atypical events in quantum information protocols. Evaluation of the probability of atypical events may be easy for independent and identically distributed (i.i.d.) quantum states. There, one can apply Chernoff-Hoeffding theorem [1] or Sanov’s theorem [2, 3] to obtain an upper bound on the deviation from the expectation value for an additive observable allowing a small failure probability. For certain applications, however, one even needs a concentration inequality that applies to an adversarial setup. This includes quantum key distribution (QKD), verifiable blind quantum computation [4], and verification and learning of a quantum state prepared by an adversary [5], where the underlying statistics may be governed by an adversary.
@@
In the field of QKD, for example, proving the security against collective attacks in a finite number of communication rounds is relatively easy, and many QKD protocols have been proved secure at least against collective attacks even in the finite-size regime. The situation gets involved when considering the security against general attacks. There are two mainstreams in the way of proving the security against general attacks in finite-size regimes. One of them is called the phase-error correction method [6, 7, 8, 9], which is the main focus of this paper, and evaluates the number of phase error patterns or a phase error rate while allowing an exponentially small failure probability. Correcting the phase error with a failure probability $\epsilon$ ensures $\sim\epsilon^{1/2}$-secrecy of the protocol against the adversary. In Ref. [10], it was shown that in the phase-error correction method, imposing a permutation invariance to the protocol either actively or passively reduces the problem of general attacks to the collective attacks at the cost of increasing the failure probability $\epsilon$ by a polynomial factor. This implies that the error exponent of the failure probability of the phase error correction is the same against collective attacks and against general attacks. After that, a distinct approach using Azuma’s inequality, which is based on sequentiality of measurements instead of permutation invariance, was proposed [11]. Since Azuma’s inequality is easier to evaluate, it prevails in the QKD community. At the cost of its simplicity, the error exponent of the failure probability when using Azuma’s inequality is always less than that of the collective attack case. Intensive research has been made on tightening Azuma’s inequality with additional information [12, 13, 14] including the recently found one with unconfirmed knowledge [14], but it is still looser than the Sanov-type inequalities [2, 3] that apply to the i.i.d. random variables. Note that in another approach to the security proofs based on the leftover hash lemma against a quantum adversary [15], a reduction of the security against general attacks to that against collective attacks by exploiting the permutation symmetry has also been studied under the name of de Finetti reduction and the post-selection technique [16, 17, 18, 19, 15, 20, 21]. The problems with these approaches were that the reduction was not very straightforward or that they led to worse performance. Then, the community gradually shifted to an alternative method that uses more recently developed entropy accumulation theorem [22, 23, 24]. Interestingly, both Azuma’s inequality and the entropy accumulation utilize sequential conditionings by previous measurement outcomes, yet they may be looser than the bound that applies to the i.i.d. case.
@@
In this paper, we revisit the idea based on permutation invariance and derive concentration inequalities through the reduction to the i.i.d. scenario. The first case we consider is when a quantum state on a multi-qudit system is invariant under the permutation of the subsystems possibly with additional symmetry on each subsystem. We show that whenever an outcome of the quantum measurement has a small appearance probability for any i.i.d. quantum state, it also has a small appearance probability for these permutation-invariant states up to an explicit factor polynomial in the number of subsystems. This can be proved by observing that any permutation-invariant quantum state is bounded from above by the mixture of i.i.d. quantum states up to a polynomial factor, i.e., we have the following.
Theorem 1 (Informal).
Let $\rho_{A^{n}}$ be a permutation-invariant state on ${\cal H}_{A}^{\otimes n}$ with $\dim{\cal H}_{A}=d$. Then, we have
[…]
where $f_{q}(n,d)={\cal O}(n^{(d^{2}-1)/2})$ for a large $n$.
Compared to the known results [15, 20], we tightly evaluate the polynomial prefactor and improve the bound for a general permutation-invariant state. We also obtain a bound that explicitly depends on the irreducible representation space of the permutation group. In particular, when restricted to the totally symmetric subspace, our bound reproduces the one developed for the post-selection technique [15, 20]. The second case we consider is when we perform independent and identical quantum measurements on an arbitrary quantum state prepared by an adversary. In this case, we have an inequality between the probability mass function for an empirical probability of the measurement outcomes and a mixture of multinomial distribution up to a polynomial factor. We then obtain a Sanov-type concentration inequality for the probability of obtaining an empirical probability that lies in a set ${\cal A}$ of atypical events.
Theorem 3 (Informal).
Let ${\cal M}$ be a $k$-outcome measurement channel and ${\cal P}$ be the probability $(k-1)$-simplex. For a state $\rho$, ${\cal M}(\rho)$ corresponds to a point in ${\cal P}$. Let ${\cal R}\subset{\cal P}$ be a region that is bounded by a hyperplane and contains $\{{\cal M}(\rho):\forall\rho\}$ and all the vertices of ${\cal P}$ except one. Expand ${\cal R}$ by moving the boundary between ${\cal R}$ and ${\cal P}\setminus{\cal R}$ by ${\cal O}(n^{-1})$ to define the region ${\cal R}^{\prime}$. Let $P_{\bm{x}}$ denotes the empirical probability for a sequence $\bm{x}$. Then, for any closed convex subset ${\cal A}\subseteq{\cal P}$ and for any $n$-partite quantum state $\rho^{n}$, the sequence $\bm{x}$ of outcomes of the measurement ${\cal M}^{\otimes n}$ performed on $\rho^{n}$ satisfies
${\cal M}^{\otimes n}(\rho^{n})[P_{\bm{x}}\in{\cal A}]\leq\max_{\bm{p}\in{\cal A}}\max_{\bm{q}\in{\cal R}^{\prime}}f_{c}(n,k)e^{-nD(\bm{p}\|\bm{q})},$ (2)
where $f_{c}(n,k)={\cal O}(n^{(k-1)/2})$ for a large $n$ and $D(\cdot\|\cdot)$ denotes the Kullback-Leibler divergence.
As opposed to the first case, the polynomial factor in this case does not depend on the dimension of the underlying quantum system but depends on the number of measurement outcomes.
@@
Applicability of our results is not restricted to a finite-size analysis of QKD. In the verifiable blind quantum computation [4] and learning of a non-i.i.d. quantum state [5] as mentioned earlier, the failure probability of a randomly-permuted property test on an adversarial quantum state is estimated. Our results may fit to these cases as well, where one cares about exponentially rare outcomes of measurements performed on an adversarial state.
1.2 Summary of our main results
@@
$\begin{split}Z_{n}^{d}&\coloneqq\biggl{\{}\bm{n}=(n_{1},\ldots,n_{d})\in\{0,\ldots,n\}^{\times d}:\\ &\hskip 113.81102pt\sum_{i=1}^{d}n_{i}=n\biggr{\}},\end{split}$ (5)
@@
$\begin{split}&\mathrm{supp}({\cal S}_{A^{n}}(P_{\bm{n}}\rho_{A^{n}}P_{\bm{n}}))\\ &\leq f_{q}(n,d;\bm{n})\\ &\qquad\times{\cal S}_{A^{n}}\!\Bigl{(}\mathbb{E}_{U_{A}\sim{\rm Haar}(d)}\Bigl{[}\Bigl{(}U_{A}\,\mathrm{diag}\Bigl{(}\frac{\bm{n}}{n}\Bigr{)}\,U_{A}^{\dagger}\Bigr{)}^{\otimes n}\Bigr{]}\Bigr{)},\end{split}$ (19)
@@
where we used the fact that $\tau_{A}^{\otimes n}$ is permutation invariant and thus satisfies ${\cal R}_{A^{n}}\circ{\cal S}_{A^{n}}(\tau_{A}^{\otimes n})=\tau_{A}^{\otimes n}$. We see that Eq. (23) takes a similar form to Eq. (3) but with different coefficients $f_{q}(n,d;\bm{n})$. For comparison, let us first reformulate the right-hand side of Eq. (23) using Lemma 1 as
$\begin{split}\rho_{A^{n}}&\leq f_{q}(n,d)\sum_{\bm{n}\in Y_{n}^{d}}\mathrm{Tr}[P_{\bm{n}}\rho_{A^{n}}]\\ &\qquad\times\mathbb{E}_{U_{A}\sim{\rm Haar}(d)}\Bigl{[}\Bigl{(}U_{A}\,\mathrm{diag}\Bigl{(}\frac{\bm{n}}{n}\Bigr{)}\,U_{A}^{\dagger}\Bigr{)}^{\otimes n}\Bigr{]}.\end{split}$ (24)
Then, the right-hand side is in the form of a probabilistic mixture of i.i.d. states up to the factor $f_{q}(n,d)$. When $n\gg d$, which is of our interest, the function $f_{q}(n,d)$ scales as $n^{(d^{2}-1)/2}$ whereas the coefficient in the conventional bound (3) scales as $n^{d^{2}-1}$. Our bound thus achieves the square-root improvement in this case.
@@
$\begin{split}&\leq f_{q}(n,d;\bm{n})f_{q}(m,d^{\prime};\bm{m})\\ &\quad\times{\cal S}_{A^{n}}\!\biggl{(}\mathbb{E}_{U_{A}\sim{\rm Haar}(d)}\Bigl{[}\Bigl{(}U_{A}\,\mathrm{diag}\Bigl{(}\frac{\bm{n}}{n}\Bigr{)}\,U_{A}^{\dagger}\Bigr{)}^{\otimes n}\Bigr{]}\biggr{)}\\ &\qquad\otimes{\cal S}_{B^{m}}\!\biggl{(}\mathbb{E}_{U_{B}\sim{\rm Haar}(d^{\prime})}\Bigl{[}\Bigl{(}U_{B}\,\mathrm{diag}\Bigl{(}\frac{\bm{m}}{m}\Bigr{)}\,U_{B}^{\dagger}\Bigr{)}^{\otimes m}\Bigr{]}\biggr{)},\end{split}$ (29)
@@
${\cal H}_{\bm{j}_{0}(\bm{m})}^{R}\otimes{\cal H}_{\bm{j}_{0}(\bm{m})}^{D}$
$=\bigotimes_{j=1}^{k}\bigl{(}{\cal H}_{j}^{R}\otimes{\cal H}_{j}^{D}\bigr{)}^{\otimes n_{j}}$ (41)
@@
${\cal R}^{\prime}\coloneqq\left\{\bm{p}\in{\cal P}:\bm{u}\cdot\bm{p}\leq\frac{\sqrt{2}\|\bm{u}-(\bm{v}\cdot\bm{u})\bm{v}\|}{n}\right\},$ (58)
@@
Consider the measurement ${\cal M}^{\otimes n}$ on $n$ systems $A^{n}$, which has the POVM $\{M^{\otimes n}(\bm{i})\}_{\bm{i}\in{\cal X}^{\times n}}$ with $M^{\otimes n}\bigl{(}(i_{1},\ldots,i_{n})\bigr{)}=M(i_{1})\otimes\cdots\otimes M(i_{n})$. Let ${\cal A}\subseteq{\cal P}$ be a non-empty closed convex region of the probability simplex. Then, for any quantum state $\rho\in{\cal D}({\cal H}^{\otimes n})$, the probability mass function ${\cal M}^{\otimes n}(\rho):\bm{i}\mapsto\mathrm{Tr}[M^{\otimes n}(\bm{i})\rho]$ satisfies
$\begin{split}&\sum_{\bm{i}:\bm{\chi}(\bm{i})/n\in{\cal A}}{\cal M}^{\otimes n}(\rho)(\bm{i})\\ &\leq\max_{\bm{p}\in{\cal A},\bm{q}\in{\cal R}^{\prime}}f_{c}(n,k)\exp[-nD(\bm{p}\|\bm{q})],\end{split}$ (59)
@@
Figure 1: Geometrical relations between the (convex) subsets ${\cal Q}$, ${\cal R}$, ${\cal R}^{\prime}$, and ${\cal A}$ of ${\cal P}$. The subset ${\cal A}$ is non-empty, closed, and convex. The subset ${\cal R}$ is bounded by a hyperplane and contains ${\cal Q}$ and all the extremal points of ${\cal P}$ except $\bm{p}_{1}=(1,0,\ldots,0)$. The set ${\cal R}^{\prime}$ consists of every point whose distance to ${\cal R}$ is no larger than $\sqrt{2}/n$.
@@
1.3 Organization of the paper
In Sec. 2, we prove Theorems 13 stated above, where each subsection is assigned for each theorem. In Sec. 3, we apply our results to simple examples and numerically demonstrate the tightness of our bounds by comparison with conventionally used bounds. The first example in Sec. 3.1 is an estimation task of measurement outcomes with two non-orthogonal projections, which is not directly related to real applications but is still informative. The second example is an application to the quantum key distribution protocol. Finally, in Sec. 4, we summarize our results and discuss possible applications.
2 Proofs of the main theorems
2.1 A permutation-invariant quantum state is bounded by a mixture of i.i.d. quantum states
@@ 2 Proofs of the main theorems
$\begin{split}{\cal S}_{A^{n}}(\tau_{A^{n}})=\bigoplus_{\bm{n}\in Y_{n}^{d}}\mathrm{Tr}^{(\bm{n})}_{{\cal V}_{\bm{n}}}(P_{\bm{n}}\tau_{A^{n}}P_{\bm{n}}|_{{\cal U}_{\bm{n}}\otimes{\cal V}_{\bm{n}}})\otimes\ket{0}\!\bra{0}_{{\cal V}_{\bm{n}}},\end{split}$ (61)
where $\mathrm{Tr}^{(\bm{n})}$ denotes the trace acting on a matrix on ${\cal U}_{\bm{n}}\otimes{\cal V}_{\bm{n}}$, $\mathrm{Tr}^{(\bm{n})}_{{\cal V}_{\bm{n}}}$ denotes the partial trace over the space ${\cal V}_{\bm{n}}$, and |$\ket{0}\!\bra{0}_{{\cal V}_{\bm{n}}}$ is an arbitrary pure state on ${\cal V}_{\bm{n}}$. We also define another CPTP map ${\cal R}_{A^{n}}$ as follows:
$\begin{split}{\cal R}_{A^{n}}(\tau_{A^{n}})=\bigoplus_{\bm{n}\in Y_{n}^{d}}\mathrm{Tr}^{(\bm{n})}_{{\cal V}_{\bm{n}}}(P_{\bm{n}}\tau_{A^{n}}P_{\bm{n}}|_{{\cal U}_{\bm{n}}\otimes{\cal V}_{\bm{n}}})\otimes\frac{I_{{\cal V}_{\bm{n}}}}{\dim{\cal V}_{\bm{n}}}.\end{split}$ (62)
@@
$\begin{split}&\mathrm{supp}({\cal S}_{A^{n}}(P_{\bm{n}}\rho_{A^{n}}P_{\bm{n}}))\\ &=\begin{cases}\mathrm{supp}(\rho_{{\cal U}_{\bm{n}}})\otimes\ket{0}\!\bra{0}_{{\cal V}_{\bm{n}}}&\text{on }{\cal U}_{\bm{n}}\otimes{\cal V}_{\bm{n}}\\ 0&\text{otherwise}\end{cases}.\end{split}$
@@
$\begin{split}&=\frac{1}{\alpha_{\bm{n}}(\bm{t})\dim{\cal V}_{\bm{n}}}\\ &\quad\times{\cal S}_{A^{n}}(P_{\bm{n}}\mathbb{E}_{U_{A}\sim\text{Haar}(d)}[(U_{A}\mathrm{diag}(\bm{t})U_{A}^{\dagger})^{\otimes n}]P_{\bm{n}})\end{split}$ (71)
$\begin{split}&\leq\frac{\dim{\cal U}_{\bm{n}}}{s_{\bm{n}}(\bm{t})\dim{\cal V}_{\bm{n}}}\\ &\quad\times{\cal S}_{A^{n}}(\mathbb{E}_{U_{A}\sim\text{Haar}(d)}[(U_{A}\mathrm{diag}(\bm{t})U_{A}^{\dagger})^{\otimes n}]).\end{split}$ (72)
@@
$\begin{split}&\mathrm{supp}({\cal S}_{A^{n}}(P_{\bm{n}}\rho_{A^{n}}P_{\bm{n}}))\\ &\leq\min_{\bm{t}=(t_{1},\ldots,t_{d})\geq 0:\sum_{i=1}^{d}t_{i}=1}\frac{\dim{\cal U}_{\bm{n}}}{s_{\bm{n}}(\bm{t})\dim{\cal V}_{\bm{n}}}\\ &\qquad\times{\cal S}_{A^{n}}\!\Bigl{(}\mathbb{E}_{U_{A}\sim{\rm Haar}(d)}\Bigl{[}\bigl{(}U_{A}\,\mathrm{diag}(\bm{t})\,U_{A}^{\dagger}\bigr{)}^{\otimes n}\Bigr{]}\Bigr{)}\end{split}$ (75)
@@
2.2 Generalization to the case with symmetry restrictions